An irregularly-shaped iceberg has a volume of 3580 m3. It floats with 90% of its volume

underwater, in sea water with a density of 1025 kg/m3. The bottom of the iceberg hangs

18.5 m below the surface of the water.

a. What is the absolute pressure in the water at the bottom edge of the iceberg?

b. What is the density of the iceberg?

c. What is the mass of the iceberg?
Answer #1

(a) Pressure, Pabs = dwater*g*h + Po

Pabs = 1025*9.8*18.5 + 101325 = 185832.5 Pa + 101325 Pa

Pabs = 287157.5 Pa

(b) As the iceberg floats on sea water, the ratio of density incase of floatation is given as,

dice/dwater = Vimmersed/Vice

dice = 1025*90/100 = 922.5 Kg/m3

(c) Mass of iceberg, mice = dice*Volume = 922.5*3580

mice = 3302550 Kg = 3.3*106 Kg

I hope i was able to answer your question. If you need anymore explanation, please ask in comments. I would be happy to help.

Please rate the solution if you are satisfied with the solution.